OpenStudy (anonymous):

consider the following quadratic function...k(x)=(x-6)..find the vertex of this function

OpenStudy (anonymous):

that's not a quadratic function

OpenStudy (anonymous):

is (x-6) raised to the 2nd power?

OpenStudy (anonymous):

no...it's not on the problem...that's how it's worded

OpenStudy (anonymous):

k(x)=x-6 is a line

OpenStudy (anonymous):

I have the answer..I just don't understand how to do the problem...the answer is (6,0)

OpenStudy (anonymous):

Well, I'm going to assume it is k(x)=(x-6)^2 Expand: x^2-12x+36 The vertex of a parabola written in the for ax^2+bx+c is: \[(-\frac{b}{2a},f(-\frac{b}{2a})\] \[(\frac{12}{2},f(\frac{12}{2}))=(6,f(6))=(6,0)\]

OpenStudy (anonymous):

looks good....worked it out..now quick question...why is f(6)=0?

OpenStudy (anonymous):

f(6)=6^2-12(6)+36=36-72+36=0

OpenStudy (anonymous):

or in the original form, f(x)=(x-6)^2 f(6)=(6-6)^2=0^2=0

OpenStudy (anonymous):

makes sense now thank you